# Tuan Duc Ngo UMass Amherst &Ashkan Mirzaei Snap Inc. &Guocheng Qian

### v1

@@ Abstract
We propose a novel algorithm for accelerating dense long-term 3D point tracking in videos. Through analysis of existing state-of-the-art methods, we identify two major computational bottlenecks. First, transformer-based iterative tracking becomes expensive when handling a large number of trajectories. To address this, we introduce a coarse-to-fine strategy that begins tracking with a small subset of points and progressively expands the set of tracked trajectories. The newly added trajectories are initialized using a learnable interpolation module, which is trained end-to-end alongside the tracking network. Second, we propose an optimization that significantly reduces the cost of correlation feature computation, another key bottleneck in prior methods. Together, these improvements lead to a 5–100$\times$ speedup over existing approaches while maintaining state-of-the-art tracking accuracy.
Figure 1: DELTAv2 achieves state-of-the-art for dense 3D tracking — matching the accuracy of DELTA while being 5×5$\times$5 × faster, and outperforming all prior methods in speed–accuracy tradeoff. (Left) Long-range 3D trajectories on real-world videos. (Right) Performance vs. FPS comparison.
@@ 1 Introduction
Together, these optimizations achieve a 5×5$\times$5 × speed-up over DELTA on dense 3D tracking of 100-frame videos, while maintaining state-of-the-art accuracy (Fig. 1).
In summary, our main contributions are: 1) a coarse-to-fine tracking algorithm that starts with sparsely sampled trajectories and densifies them over iterations; 2) a learnable interpolator that produces dense trajectories in each iteration and supports adaptive resampling; 3) an accelerated 4D correlation computation, bringing dense 3D tracking closer to real-time.
@@ 2 Related Work
Downsample video reso. (4×4$\times$4 ×) 35.8 383
Subsample video frames (4×4$\times$4 ×) 65.4 833
Subsample trajectories (16×16$\times$16 ×) 71.2 585
@@
Optical Flow estimates motion by establishing dense correspondences between consecutive frames. Classical variational methods memin1998dense ; horn1980determining ; brox2004high faced limitations in handling complex dynamics such as fast motion, occlusions, and large displacements. The advent of CNN-based techniques (flownet, ; ranjan2017optical, ; xu2017accurate, ; pwcnet, ) brought substantial improvements to short-term motion estimation. RAFT (raft, ) introduced a key innovation by computing dense correlation volumes across all pixel pairs. Later approaches expanded on this architecture by introducing transformer-based tokenization of correlation volumes (flowformer, ), employing global feature aggregation to better address occlusions (jiang2021learning, ), and modeling optical flow as a softmax-based matching task (gmflow, ). Attempts to extend optical flow to long-term sequences using multi-frame methods (raft, ; godet2021starflow, ; shi2023videoflow, ) or point-tracking integration (dot, ; flowtrack, ) often face challenges with drift and occlusion, limiting long-term tracking performance.
@@
Scene Flow extends optical flow into 3D by estimating dense motion in 3D space. Some methods rely on RGB-D inputs (hadfield2011kinecting, ; hornacek2014sphereflow, ; quiroga2014dense, ; raft3d, ; yang2021learning, ), while others operate on monocular video alone (hur2020self, ; hur2021self, ; bayramli2023raft, ; jiang2023emr, ). A separate line of work estimates 3D motion directly from point clouds (flownet3d, ; flownet3d++, ; hplflownet, ; niemeyer2019occupancy, ). Several approaches (he2022learning, ; liu2019meteornet, ; huang2022dynamic, ) leverage temporal context to predict scene flow from videos. More recently, methods such as (sucar2025dynamic, ; liang2025zero, ) build on top of foundation 3D reconstruction models (dust3r, ; monst3r, ; mast3r_arxiv24, ) to estimate 3D scene flow in a more general and scalable manner.
@@
Point Tracking estimates long-range motion trajectories in videos. Particle Video (particle, ) introduced particle trajectories for long-range video motion. TAP-Vid (tapvid, ) provided a comprehensive benchmark to evaluate point tracking and TAPNet, a baseline that predicts tracking locations using global cost volume and soft-argmax operation. PIPs (pips, ) revisited the concept of particle video and proposed a network that updates trajectories iteratively over fixed temporal windows, but ignored spatial context with independent point tracking and struggle with occlusion. TAPIR (tapir, ) combines the global estimation of TAPNet (tapvid, ) and the iterative refinement of (pips, ) with variable-length window. Subsequent efforts further improve tracking performance by jointly process multiple points with temporal attention and spatial attention (cotracker, ); using 4D correlation feature (cho2024local, ). CoTracker3 (cotracker3, ) proposes a lightweight architecture and a semi-supervised training strategy to further improve the 2D tracking performance. SceneTracker (scenetracker, ) and SpaTracker (spatialtracker, ) extend point tracking to 3D by incorporating depth information, but remain inefficient for dense tracking due to computationally expensive cross-track attention. DELTA (ngo2024delta, ) is the first approach tackling the dense 3D tracking problem with an efficient transformer and upsample layer to obtain high-resolution dense tracking. Concurrently, several other efforts also explore 3D tracking. Seurat (seurat, ) estimates depth changes on top of a 2D tracker (cotracker, ; cho2024local, ) to recover 3D motion. TAPIP3D (tapip3d, ) proposes a 3D-space correlation mechanism for directly tracking points in 3D space. St4RTrack (st4rtrack, ) leverages recent 3D reconstruction models (dust3r, ; monst3r, ) and extends them for dense 3D tracking via joint optimization.
@@
Tracking by Reconstructing tackles long-range motion estimation by explicitly reconstructing a deformable scene representation. OmniMotion (wang2023tracking, ) jointly optimizes a NeRF (mildenhall2020nerf, ) with a bijective deformation field (Dinh2016DensityEU, ) and extracts 2D point trajectories through this mapping. More recent efforts leverage DINOv2 (oquab2023dinov2, ) to compute long-range correspondences, either through enhanced invertible deformation fields (Song2024TrackEE, ) or self-supervised techniques (dino_tracker_2024, ). Shape-of-Motion (wang2024shape, ) uses 3D Gaussian Splatting to jointly learn geometry and motion, enabling point tracking by tracing Gaussian positions across frames. While these reconstruction-based methods are capable of producing dense trajectories, they rely on computationally intensive per-video optimization. As a result, they remain significantly slower and less accurate than feedforward, data-driven tracking models on standard benchmarks.
@@ 3 Method
[…]
@@ 3.1 Preliminary: Modern Tracking Frameworks
Our work builds upon a modern tracking framework. We select DELTA (ngo2024delta, ), a recent approach for dense 3D tracking from RGB-D videos, but other frameworks have a similar structure (See Fig. 3 for an overview). First, a feature extraction network $F$ (similar to raft ) creates image feature map for each frame. Then, the core of the algorithm iteratively evolves a data structure of size of dense tracks $\mathcal{D}\in\mathbb{R}^{T\times\frac{H}{r}\times\frac{W}{r}}$, denoted as $\{{P}_{i}\}$, where $i$ is the trajectory index, and ${P}_{i}=[{\bm{p}}_{1}^{i},{\bm{p}}_{2}^{i},\cdots,{\bm{p}}_{T}^{i}]$, with ${\bm{p}}_{t}^{i}=(u_{t}^{i},v_{t}^{i},d_{t}^{i},o_{t}^{i})$ being the 3D location and visibility of the point associated with the $i$-th trajectory at the $t$-th frame. The downsampling factor $r$ is beneficial for faster processing.
@@
where $\mathcal{F}_{t}^{i}$ are image features extracted by $F$, $Corr_{t}^{i}$ are correlation features (raft, ; pips, ; cotracker, ; cho2024local, ; ngo2024delta, ), $DCorr_{t}^{m}$ is depth correlation, and $\gamma_{pos}$ and $\gamma_{time}$ are the positional embedding of the input position ${\bm{x}}_{t}^{i}=(u_{t}^{i},v_{t}^{i},d_{t}^{i})$ and time $t$, respectively. For more details about the individual components, refer to ngo2024delta .
@@ 3.3.2 Learnable Interpolation Module
Let $\mathcal{F}\in\mathbb{R}^{\frac{H}{r}\times\frac{W}{r}\times C_{\mathcal{F}}}$ denote the dense feature map of the query frame of the dense tracking from the feature extractor $F$. The initial query feature $\mathcal{F}_{(u,v)}$ is first refined via $L$ multi-head cross-attention blocks over the support set. We follow the Alibi attention scheme (alibi, ) in these blocks, where a relative positional bias is added to the attention logits, computed as the L1 distance between the query and support pixel locations. The refined query and the final weights are predicted as:
~$\tilde{\mathcal{F}}_{(u,v)}=\text{CrossAttn}^{(L)}\left(\mathcal{F}_{(u,v)},\left\{\mathcal{F}_{(u_{j},v_{j})}\right\}\right),\>\mathbf{w}_{(u,v)}=\text{softmax}\left({\bm{q}}(\tilde{\mathcal{F}}_{(u,v)})\cdot{\bm{k}}\left(\left\{\mathcal{F}_{(u_{j},v_{j})}\right\}\right)\right)$ (3)
@@ 3.3.3 Accelerating Computation of Correlation Features
The main inefficiency arises from the input channel size (typically 7×7=497$\times$ 7=497 × 7 = 49), which is not divisible by 8, leading to poor GPU utilization in dense settings. To address this, we add a lightweight MLP projection that reduces the dimension to 32, followed by LayerNorm and ReLU. This adds minimal overhead for sparse tracking and yields a significant speedup in the dense setting. An alternative from cotracker3 replaces the dual-conv module with pure MLPs on flattened 4D correlation, but we find it performs worse on dense 2D/3D tracking (See Tab. 6).
@@ 4 Experiments
Table 2: Long-range optical flow results on CVO (accflow, ; dot, ).
@@
Implementation Details: Following prior work (ngo2024delta, ; cotracker, ; spatialtracker, ), we use the Kubric engine (kubric, ) to generate 5,631 RGB-D training videos with both sparse and dense tracking annotations. During training, we apply online data augmentation, including color jittering, random scaling, and random cropping. As in ngo2024delta , we supervise the model on both sparse and dense tracks. The total loss combines the 2D coordinate loss $\mathcal{L}_{2D}$, inverse depth loss $\mathcal{L}_{depth}$, and visibility loss $\mathcal{L}_{visib}$. The loss weights $\lambda_{2d},\lambda_{depth},\lambda_{visib}$ follow those in ngo2024delta . We initialize the model from the pretrained DELTA checkpoint and add parameters for the interpolation module. The model is trained for 100,000 iterations on 8 A100 GPUs using AdamW with a one-cycle learning rate schedule, starting from an initial learning rate of $2\times 10^{-4}$.
@@ 4.1 Comparison with prior work
Benchmark Datasets. We evaluate our method on a diverse set of tracking benchmarks covering 2D optical flow, dense 3D pixel tracking, and sparse 3D point tracking. (1) Long-range 2D optical flow: We use the CVO dataset (accflow, ; dot, ), which includes three splits: CVO-Clean, CVO-Final, and CVO-Extended. Each split contains about 500 videos with 7 frames (48 frames for CVO-Extended), captured at 60 FPS, with dense long-range optical flow annotations and occlusion masks. (2) Dense 3D pixel tracking: Following ngo2024delta , we evaluate on a held-out test split of the Kubric dataset (kubric, ), which includes 143 RGB-D videos, each 24 frames long, with dense ground-truth 3D trajectories for every pixel. (3) 3D point tracking: We benchmark on the large-scale TAP-Vid3D dataset (tapvid3d, ), which contains 4,569 videos from DriveTrack (balasingam2024drivetrack, ), PStudio (joo2017panoptic, ), and Aria (pan2023aria, ), covering a range of real-world and simulated environments, with video lengths from 25 to 300 frames.
@@
Dense 3D Tracking. Table 3 reports results on the Kubric3D test set. Our method matches the performance of DELTA while significantly outperforming other baselines, including 3D tracking methods such as SpaTrackerspatialtracker and SceneTrackerscenetracker , as well as 2D-to-3D lifted approaches like CoTracker2cotracker and CoTracker3cotracker3 . Importantly, our approach is substantially more efficient: it achieves up to a 5×5$\times$5 × speed-up over DELTA and is approximately 100×100$\times$100 × faster than SpaTracker, enabling practical deployment in real-time or large-scale applications. Qualitative results of dense 3D tracking on in-the-wild videos are provided in the supplementary materials.
@@ 4.2 Ablation Study
We conduct a series of ablation studies on the CVO Extended split dot . Runtime is measured as the average time required to densely track all 384×512384$\times$ 512384 × 512 pixels across a 48-frame video using a machine with a single A100 GPU. The number of iterations is set to 4 by default, unless otherwise specified.
Analysis of the Coarse-to-fine strategy. Figure 5a compares our coarse-to-fine strategy (schedule: 8,4,2,1) with the baseline that tracks all trajectories in every iteration. Our method achieves similar accuracy while being 2.75×2.75$\times$2.75 × faster. In Fig.5b, we ablate different interpolators, showing that our attention-based design outperforms fixed alternatives like nearest-neighbor and bilinear. Fig. 5c compares different subsampling schedules, where $(s_{1},s_{2},s_{3},s_{4})$ denotes the subsampling factors across 4 iterations. A finer schedule, such as (2,2,2,1), improves accuracy but with increased runtime.
@@
Runtime Analysis. Table 6 summarizes our runtime improvements. Replacing the original 4D correlation in DELTA ngo2024delta reduces runtime by 33%. Adding our coarse-to-fine strategy yields a further 3×3$\times$3 × speed-up. A 3-layer transformer brings runtime to 5.8s, though with some loss in accuracy.

### v2

# DELTAv2: Accelerating Dense 3D Tracking

@@ Abstract
[…]
Figure 1: DELTAv2 achieves state-of-the-art for dense 3D tracking — matching the accuracy of DELTA while being $5\times$ faster, and outperforming all prior methods in speed–accuracy tradeoff. (Left) Long-range 3D trajectories on real-world videos. (Right) Performance vs. FPS comparison.
@@ 1 Introduction
Together, these optimizations achieve a $5\times$ speed-up over DELTA on dense 3D tracking of 100-frame videos, while maintaining state-of-the-art accuracy (Fig. 1).
In summary, our main contributions are: 1) a coarse‑to‑fine tracking algorithm that starts with sparsely sampled trajectories and densifies them over iterations; 2) a learnable interpolator that produces dense trajectories in each iteration and supports adaptive resampling; 3) an accelerated 4D correlation computation, bringing dense 3D tracking closer to real‑time.
@@ 2 Related Work
Downsample video reso. ($4\times$) 35.8 383
Subsample video frames ($4\times$) 65.4 833
Subsample trajectories ($16\times$) 71.2 585
@@
Optical Flow estimates motion by establishing dense correspondences between consecutive frames. Classical variational methods memin1998dense; horn1980determining; brox2004high faced limitations in handling complex dynamics such as fast motion, occlusions, and large displacements. The advent of CNN-based techniques (flownet; ranjan2017optical; xu2017accurate; pwcnet) brought substantial improvements to short-term motion estimation. RAFT (raft) introduced a key innovation by computing dense correlation volumes across all pixel pairs. Later approaches expanded on this architecture by introducing transformer-based tokenization of correlation volumes (flowformer), employing global feature aggregation to better address occlusions (jiang2021learning), and modeling optical flow as a softmax-based matching task (gmflow). Attempts to extend optical flow to long-term sequences using multi-frame methods (raft; godet2021starflow; shi2023videoflow) or point-tracking integration (dot; flowtrack) often face challenges with drift and occlusion, limiting long-term tracking performance.
@@
Scene Flow extends optical flow into 3D by estimating dense motion in 3D space. Some methods rely on RGB-D inputs (hadfield2011kinecting; hornacek2014sphereflow; quiroga2014dense; raft3d; yang2021learning), while others operate on monocular video alone (hur2020self; hur2021self; bayramli2023raft; jiang2023emr). A separate line of work estimates 3D motion directly from point clouds (flownet3d; flownet3d++; hplflownet; niemeyer2019occupancy). Several approaches (he2022learning; liu2019meteornet; huang2022dynamic) leverage temporal context to predict scene flow from videos. More recently, methods such as (sucar2025dynamic; liang2025zero) build on top of foundation 3D reconstruction models (dust3r; monst3r; mast3r_arxiv24) to estimate 3D scene flow in a more general and scalable manner.
@@
Point Tracking estimates long-range motion trajectories in videos. Particle Video (particle) introduced particle trajectories for long-range video motion. TAP-Vid (tapvid) provided a comprehensive benchmark to evaluate point tracking and TAPNet, a baseline that predicts tracking locations using global cost volume and soft-argmax operation. PIPs (pips) revisited the concept of particle video and proposed a network that updates trajectories iteratively over fixed temporal windows, but ignored spatial context with independent point tracking and struggle with occlusion. TAPIR (tapir) combines the global estimation of TAPNet (tapvid) and the iterative refinement of (pips) with variable-length window. Subsequent efforts further improve tracking performance by jointly process multiple points with temporal attention and spatial attention (cotracker); using 4D correlation feature (cho2024local). CoTracker3 (cotracker3) proposes a lightweight architecture and a semi-supervised training strategy to further improve the 2D tracking performance. SceneTracker (scenetracker) and SpaTracker (spatialtracker) extend point tracking to 3D by incorporating depth information, but remain inefficient for dense tracking due to computationally expensive cross-track attention. DELTA (ngo2024delta) is the first approach tackling the dense 3D tracking problem with an efficient transformer and upsample layer to obtain high-resolution dense tracking. Concurrently, several other efforts also explore 3D tracking. Seurat (seurat) estimates depth changes on top of a 2D tracker (cotracker; cho2024local) to recover 3D motion. TAPIP3D (tapip3d) proposes a 3D-space correlation mechanism for directly tracking points in 3D space. St4RTrack (st4rtrack) leverages recent 3D reconstruction models (dust3r; monst3r) and extends them for dense 3D tracking via joint optimization.
@@
Tracking by Reconstructing tackles long-range motion estimation by explicitly reconstructing a deformable scene representation. OmniMotion (wang2023tracking) jointly optimizes a NeRF (mildenhall2020nerf) with a bijective deformation field (Dinh2016DensityEU) and extracts 2D point trajectories through this mapping. More recent efforts leverage DINOv2 (oquab2023dinov2) to compute long-range correspondences, either through enhanced invertible deformation fields (Song2024TrackEE) or self-supervised techniques (dino_tracker_2024). Shape-of-Motion (wang2024shape) uses 3D Gaussian Splatting to jointly learn geometry and motion, enabling point tracking by tracing Gaussian positions across frames. While these reconstruction-based methods are capable of producing dense trajectories, they rely on computationally intensive per-video optimization. As a result, they remain significantly slower and less accurate than feedforward, data-driven tracking models on standard benchmarks.
@@ 3 Method
Problem setup: The input to our method is an RGB-D video, where the RGB frames are represented as $V\in\mathbb{R}^{T\times H\times W\times 3}$, with $T$, $H$, and $W$ denoting the temporal and spatial dimensions. The corresponding depth maps $D\in\mathbb{R}^{T\times H\times W}$ are generated using a monocular depth estimation model. Our model outputs dense, occlusion-aware 3D trajectories $P\in\mathbb{R}^{T\times H\times W\times 4}$, where each vector $\mathbf{p}_{(t,u,v)}=(u_{t},v_{t},d_{t},o_{t})$ captures the motion of a pixel located at $(u,v)$ in the first frame as it propagates to frame $t$. Here, $(u_{t},v_{t})$ denote the projected 2D pixel coordinates in frame $t$, $d_{t}$ is the predicted depth, and $o_{t}\in\{0,1\}$ indicates the visibility status.
@@ 3.1 Preliminary: Modern Tracking Frameworks
Our work builds upon a modern tracking framework. We select DELTA (ngo2024delta), a recent approach for dense 3D tracking from RGB-D videos, but other frameworks have a similar structure (See Fig. 3 for an overview). First, a feature extraction network $F$ (similar to raft) creates image feature map for each frame. Then, the core of the algorithm iteratively evolves a data structure of size of dense tracks $\mathcal{D}\in\mathbb{R}^{T\times\frac{H}{r}\times\frac{W}{r}}$, denoted as $\{{P}_{i}\}$, where $i$ is the trajectory index, and ${P}_{i}=[{\bm{p}}_{1}^{i},{\bm{p}}_{2}^{i},\cdots,{\bm{p}}_{T}^{i}]$, with ${\bm{p}}_{t}^{i}=(u_{t}^{i},v_{t}^{i},d_{t}^{i},o_{t}^{i})$ being the 3D location and visibility of the point associated with the $i$-th trajectory at the $t$-th frame. The downsampling factor $r$ is beneficial for faster processing.
@@
where $\mathcal{F}_{t}^{i}$ are image features extracted by $F$, $Corr_{t}^{i}$ are correlation features (raft; pips; cotracker; cho2024local; ngo2024delta), $DCorr_{t}^{m}$ is depth correlation, and $\gamma_{pos}$ and $\gamma_{time}$ are the positional embedding of the input position ${\bm{x}}_{t}^{i}=(u_{t}^{i},v_{t}^{i},d_{t}^{i})$ and time $t$, respectively. For more details about the individual components, refer to ngo2024delta.
@@ 3.3.2 Learnable Interpolation Module
Let $\mathcal{F}\in\mathbb{R}^{\frac{H}{r}\times\frac{W}{r}\times C_{\mathcal{F}}}$ denote the dense feature map of the query frame of the dense tracking from the feature extractor $F$. The initial query feature $\mathcal{F}_{(u,v)}$ is first refined via $L$ multi-head cross-attention blocks over the support set. We follow the Alibi attention scheme (alibi) in these blocks, where a relative positional bias is added to the attention logits, computed as the L1 distance between the query and support pixel locations. The refined query and the final weights are predicted as:
$\tilde{\mathcal{F}}_{(u,v)}=\text{CrossAttn}^{(L)}\left(\mathcal{F}_{(u,v)},\left\{\mathcal{F}_{(u_{j},v_{j})}\right\}\right),\>\mathbf{w}_{(u,v)}=\text{softmax}\left({\bm{q}}(\tilde{\mathcal{F}}_{(u,v)})\cdot{\bm{k}}\left(\left\{\mathcal{F}_{(u_{j},v_{j})}\right\}\right)\right)$ (3)
@@ 3.3.3 Accelerating Computation of Correlation Features
The main inefficiency arises from the input channel size (typically $7\times 7=49$), which is not divisible by 8, leading to poor GPU utilization in dense settings. To address this, we add a lightweight MLP projection that reduces the dimension to 32, followed by LayerNorm and ReLU. This adds minimal overhead for sparse tracking and yields a significant speedup in the dense setting. An alternative from cotracker3 replaces the dual-conv module with pure MLPs on flattened 4D correlation, but we find it performs worse on dense 2D/3D tracking (See Tab. 6).
@@ 4 Experiments
Table 2: Long-range optical flow results on CVO (accflow; dot).
@@
Implementation Details: Following prior work (ngo2024delta; cotracker; spatialtracker), we use the Kubric engine (kubric) to generate 5,631 RGB-D training videos with both sparse and dense tracking annotations. During training, we apply online data augmentation, including color jittering, random scaling, and random cropping. As in ngo2024delta, we supervise the model on both sparse and dense tracks. The total loss combines the 2D coordinate loss $\mathcal{L}_{2D}$, inverse depth loss $\mathcal{L}_{depth}$, and visibility loss $\mathcal{L}_{visib}$. The loss weights $\lambda_{2d},\lambda_{depth},\lambda_{visib}$ follow those in ngo2024delta. We initialize the model from the pretrained DELTA checkpoint and add parameters for the interpolation module. The model is trained for 100,000 iterations on 8 A100 GPUs using AdamW with a one-cycle learning rate schedule, starting from an initial learning rate of $2\times 10^{-4}$.
@@ 4.1 Comparison with prior work
Benchmark Datasets. We evaluate our method on a diverse set of tracking benchmarks covering 2D optical flow, dense 3D pixel tracking, and sparse 3D point tracking. (1) Long-range 2D optical flow: We use the CVO dataset (accflow; dot), which includes three splits: CVO-Clean, CVO-Final, and CVO-Extended. Each split contains about 500 videos with 7 frames (48 frames for CVO-Extended), captured at 60 FPS, with dense long-range optical flow annotations and occlusion masks. (2) Dense 3D pixel tracking: Following ngo2024delta, we evaluate on a held-out test split of the Kubric dataset (kubric), which includes 143 RGB-D videos, each 24 frames long, with dense ground-truth 3D trajectories for every pixel. (3) 3D point tracking: We benchmark on the large-scale TAP-Vid3D dataset (tapvid3d), which contains 4,569 videos from DriveTrack (balasingam2024drivetrack), PStudio (joo2017panoptic), and Aria (pan2023aria), covering a range of real-world and simulated environments, with video lengths from 25 to 300 frames.
@@
Dense 3D Tracking. Table 3 reports results on the Kubric3D test set. Our method matches the performance of DELTA while significantly outperforming other baselines, including 3D tracking methods such as SpaTrackerspatialtracker and SceneTrackerscenetracker, as well as 2D-to-3D lifted approaches like CoTracker2cotracker and CoTracker3cotracker3. Importantly, our approach is substantially more efficient: it achieves up to a $5\times$ speed-up over DELTA and is approximately $100\times$ faster than SpaTracker, enabling practical deployment in real-time or large-scale applications. Qualitative results of dense 3D tracking on in-the-wild videos are provided in the supplementary materials.
@@ 4.2 Ablation Study
We conduct a series of ablation studies on the CVO Extended split dot. Runtime is measured as the average time required to densely track all $384\times 512$ pixels across a 48-frame video using a machine with a single A100 GPU. The number of iterations is set to 4 by default, unless otherwise specified.
Analysis of the Coarse-to-fine strategy. Figure 5a compares our coarse-to-fine strategy (schedule: 8,4,2,1) with the baseline that tracks all trajectories in every iteration. Our method achieves similar accuracy while being $2.75\times$ faster. In Fig.5b, we ablate different interpolators, showing that our attention-based design outperforms fixed alternatives like nearest-neighbor and bilinear. Fig. 5c compares different subsampling schedules, where $(s_{1},s_{2},s_{3},s_{4})$ denotes the subsampling factors across 4 iterations. A finer schedule, such as (2,2,2,1), improves accuracy but with increased runtime.
@@
Runtime Analysis. Table 6 summarizes our runtime improvements. Replacing the original 4D correlation in DELTA ngo2024delta reduces runtime by 33%. Adding our coarse-to-fine strategy yields a further $3\times$ speed-up. A 3-layer transformer brings runtime to 5.8s, though with some loss in accuracy.